OpenStudy (anonymous):

A surveyor, standing 100 ft. from the base of a building, measures the angle of elevation to the top of the building to be 75 degrees. How accurately must the angle be measured for the percentage error in estimating the height of the building to be less than 4%?

OpenStudy (anonymous):

Right well Height = 100tan75 =373.205ft if we say that the angle of elevation is 'a' = 75 degrees so dH/da=100sec^2(a) and then from that we can say that: the fractional change in height is = dH/H = 0.04( for 4%) hence 0.04 = 100sec^2(a)da/H, where da is the change in angle and so da = 0.04H/(100sec^2(a)) da = 0.01 That's my take anyway